# Can Agents Fix Agent Issues?

### v2

@@ 5.1 Experimental Setup
Backbone LLMs. Based on the recent SWE leaderboard sweboard , state-of-the-art SE agents achieve higher fixing rate on general software issues when equipped with backbone LLMs GPT-4o GPT4o and Claude-3.5 Sonnet claude . Therefore, in our experiments, we mainly study how effective SE agents are in resolving agent issues with these two backbone LLMs.
Evaluation pipelines. We apply studied SE agents on AgentIssue-Bench and collect their generated patches for each issue resolution task. We then calculate the metrics of fault localization accuracy, plausible and correct resolution rates for each studied SE agent. To eliminate the randomness from LLMs, we repeat all experiments three times and present the average results.
@@
SE Agent LLM Plausibly resolved% Correctly resolved% Localization % Avg. File-level Function-level \$Cost Agentless GPT-4o 18.67 6.00 22.97 14.40 0.65 Claude-3.5-S 12.67 8.67 21.86 11.54 0.33 AutoCodeRover GPT-4o 12.67 4.67 17.05 11.30 0.23 Claude-3.5-S 17.33 12.67 25.61 18.67 0.05 SWE-agent GPT-4o 5.33 3.33 12.65 11.77 1.15 Claude-3.5-S 6.67 6.67 15.58 11.26 0.57
@@ 5.2 Quantitative Results
Overall resolution effectivenes. Table 2 shows the results of the studied SE agents on AgentIssue-Bench. In general, state-of-the-art SE agents can only correctly resolve a small number (i.e., 3.33% - 12.67%) of agent issues. In addition, in most cases, SE agents even fail to correctly identify the location (i.e., files or functions) for resolving the issue, e.g., file-level/function-level localization accuracy is less than 26%/19%. Such observations reveal the limited capabilities of state-of-the-art SE agents in understanding and resolving the issues in agent systems.
@@
Comparison among SE agents and backbone LLMs. As shown in Table 2, SE agents with Claude-3.5-S achieve higher resolution capabilities than with GPT-4o in terms of plausible resolution, correct resolution, and localization accuracy. In particular, AutoCodeRover with Claude-3.5-S achieves the highest resolved rate (i.e., 12.67%) and the highest localization accuracy (i.e., 25.61% at file level). Overall, we observe a larger potential of Claude-3.5-S in understanding agent issues than GPT-4o.
Figure 5 shows the unique and overlapped agent issues that are correctly resolved by each SE agent. We could observe that each SE agent can uniquely fix 2 - 4 bugs that cannot be resolved by any other SE agents. In addition, there is no agent issue that can be fixed by all SE agents. In other words, existing SE agents exhibit complementary capabilities to resolve agent issues.
@@
Category Resolved% Sub-category Resolved% Tool-related issues 3/9 (33.33%) Tool dependency issues 2/3 (66.67%) Tool implementation issues 1/3 (33.33%) Memory-related issues 1/8 (12.50%) Memory content errors 1/5 (20.00%) LLM operation issues 2/11 (18.18%) Model access misconfiguration 1/2 (50.00%) Prompt-related issues 1/2 (50.00%) Workflow issues 1/6 (16.67%) - - Utility issues 8/14 (57.14%) Utility implementation issues 2/6 (33.33%) Utility dependency issues 2/2 (100.00%) Utility configuration issues 4/6 (66.67%)
@@ 6 Limitation and Future work
While AgentIssue-Bench is representative of real-world agent issues by covering a wide range of different categories, the generality of our findings can still be restricted due to the current size of AgentIssue-Bench. In particular, we find reproducing issues in agent systems is significantly more challenging than in traditional software systems. Due to the nondeterminism of LLMs and changeable external resources (e.g., tools and LLM providers) interacted with agent systems, only a small number of agent issues (50 out of 201 issues) can be successfully reproduced. Moreover, huge manual effort (approximately 500 person-hours) are dedicated to preparing the Docker environment, configuring agent systems, and writing failure-triggering tests. In the future, we plan to continuously maintain and extend the benchmark to support future research on agent system maintenance.
@@ 7 Conclusion
In this work, we analyze 201 GitHub issues from 16 real-world agent systems and construct the first taxonomy of agent issues. We further build AgentIssue-Bench, the first reproducible benchmark of 50 high-quality agent issue resolution tasks. Experiments on state-of-the-art SE agents demonstrate their limited effectiveness in addressing agent issues (with resolution rates ranging from 3.33% to 12.67%), highlighting the unique challenges in maintaining agent systems and the pressing need for more advanced SE agents tailored to this emerging software paradigm.
@@ Appendix A Analyzed agent systems
Agent #stars #Loc Creation time License agent-squad [4] 5.4k 109,019 07/23/2024 Apache-2.0 License AGiXT [5] 3k 111,946 04/04/2023 MIT License autogen [8] 44.2k 197,969 08/18/2023 CC-BY-4.0, MIT License camel [10] 12.4k 206,152 03/17/2023 Apache-2.0 License babyagi [9] 21.4k 8,800 04/03/2023 MIT License CrewAI [13] 31.3k 171,395 10/27/2023 MIT License Lagent [21] 2.1k 13,075 08/20/2023 Apache-2.0 License MetaGPT [22] 55.4k 90,709 06/30/2023 MIT License RagaAI-Catalyst [26] 16.2k 47,252 10/01/2024 Apache-2.0 License ChatDev [11] 26.8k 40,478 11/04/2023 Apache-2.0 License gpt-engineer [18] 54.1k 17,460 04/29/2023 MIT License Pythagora [25] 1.8k 5,859 01/21/2023 Apache-2.0 License SWE-agent [28] 15.7k 63,388 04/02/2024 MIT License evo.ninja [16] 1.1k 31,862 08/18/2023 MIT License Superagent [27] 5.8k 58,602 05/10/2023 MIT License gpt-researcher [19] 21.3k 168,849 05/12/2023 Apache-2.0 License
@@ Appendix D Experiment Statistical Significance
SE Agent LLM Plausibly resolved% Correctly resolved% Localization % File-level Function-level Agentless GPT-4o 18.67 ($\pm{}2.31$) 6.00 ($\pm{}4.00$) 22.97 ($\pm{}4.74$) 14.40 ($\pm{}3.23$) Claude-3.5-S 12.67 ($\pm{}2.31$) 8.67 ($\pm{}2.31$) 21.86 ($\pm{}3.81$) 11.54 ($\pm{}1.36$) AutoCodeRover GPT-4o 12.67 ($\pm{}4.62$) 4.67 ($\pm{}4.62$) 17.05 ($\pm{}1.79$) 11.30 ($\pm{}1.03$) Claude-3.5-S 17.33 ($\pm{}4.62$) 12.67 ($\pm{}2.31$) 25.61 ($\pm{}3.18$) 18.67 ($\pm{}6.67$) SWE-agent GPT-4o 5.33 ($\pm{}4.62$) 3.33 ($\pm{}4.62$) 12.65 ($\pm{}5.06$) 11.77 ($\pm{}6.94$) Claude-3.5-S 6.67 ($\pm{}2.31$) 6.67 ($\pm{}2.31$) 15.58 ($\pm{}7.00$) 11.26 ($\pm{}1.59$)
@@ Appendix E Examples of Unresolved Issues
For the example depicted in Figure 7, the agent fails to identify the root cause of the KeyError, i.e., a conflict arising from generating multiple diffs for a single file. This issue is specific to the agent system, as it involves the handling of model outputs. However, instead of performing a deeper analysis, the agent merely prints a error message, resulting in an unsuccessful patch.

### v3

@@ 5.1 Experimental Setup
Backbone LLMs. Based on the recent SWE leaderboard sweboard , state-of-the-art SE agents achieve higher fixing rate on general software issues when equipped with backbone LLMs GPT-4o GPT4o and Claude-3.5 Sonnet claude . Therefore, in our experiments, we mainly study how effective SE agents are in resolving agent issues with these two backbone LLMs (temperature = 0).
Evaluation pipelines. We apply studied SE agents on AgentIssue-Bench and collect their generated patches for each issue resolution task. We then calculate the metrics of fault localization accuracy, plausible and correct resolution rates for each studied SE agent. To eliminate the randomness from LLMs, we repeat all experiments three times and present the average results. In particular, our major metric (average resolution rate over three runs) is essentially average pass@1 over three runs. Table 8 in Appendix G further presents the pass@1 and pass@3 over one run.
@@
SE Agent LLM Plausibly resolved% Correctly resolved% Localization % Avg. File-level Function-level \$Cost Agentless GPT-4o 12.00 3.33 27.82 12.99 0.65 Claude-3.5-S 12.00 4.00 27.35 17.50 0.33 AutoCodeRover GPT-4o 7.33 1.33 22.07 14.77 0.23 Claude-3.5-S 12.67 4.67 25.81 19.18 0.05 SWE-agent GPT-4o 0.67 0.67 11.67 4.22 1.15 Claude-3.5-S 2.00 2.00 9.52 6.78 0.57
@@ 5.2 Quantitative Results
Overall resolution effectiveness. Table 2 shows the results of the studied SE agents on AgentIssue-Bench. In general, state-of-the-art SE agents can only correctly resolve a small number (i.e., 0.67% - 4.67%) of agent issues. In addition, in most cases, SE agents even fail to correctly identify the location (i.e., files or functions) for resolving the issue, e.g., file-level/function-level localization accuracy is less than 28%/20%. Such observations reveal the limited capabilities of state-of-the-art SE agents in understanding and resolving the issues in agent systems.
@@
Comparison among SE agents and backbone LLMs. As shown in Table 2, SE agents with Claude-3.5-S achieve better effectiveness than with GPT-4o in terms of plausible resolution, correct resolution, and localization accuracy. In particular, AutoCodeRover with Claude-3.5-S achieves the highest resolution rate (i.e., 4.67%) and the highest function-level localization accuracy (i.e., 19.18%). Overall, we observe a larger potential of Claude-3.5-S in understanding agent issues than GPT-4o.
Figure 5 shows the unique and overlapped agent issues that are correctly resolved by each SE agent. An issue is counted as correctly resolved by an agent if it was solved in at least one of the three experimental runs. We could observe that each SE agent can uniquely fix 1 - 2 bugs that cannot be resolved by any other SE agents. In addition, there is no agent issue that can be fixed by all SE agents. In other words, existing SE agents exhibit complementary capabilities to resolve agent issues.
@@
Category Resolved% Sub-category Resolved% Tool-related issues 2/12 (16.67%) Tool dependency issues 2/3 (66.67%) LLM operation issues 1/11 (9.09%) Prompt-related issues 1/2 (50.00%) Utility issues 2/12 (16.67%) Utility configuration issues 2/6 (33.33%)
@@ 6 Limitation and Future work
While AgentIssue-Bench is representative of real-world agent issues by covering a wide range of different categories, the generality of our findings can still be restricted due to the data source and the benchmark scale. First, our work only focuses on reactive agent issues (i.e., first-expose-then-fix), as our data source (i.e., user-reported Github issues) inherently captures problems reported by agent users. This scope intentionally excludes other maintenance aspects such as preventative strategies (e.g., proactive LLM monitoring) and performance optimization, which are typically observed from the developers’ perspective using internal logs. Second, the current benchmark size is limited as reproducing issues in agent systems is significantly more challenging than in traditional software systems. Due to the nondeterminism of LLMs and changeable external resources (e.g., tools and LLM providers) interacted with agent systems, only a small number of agent issues (50 out of 201 issues) can be successfully reproduced. Moreover, huge manual effort (approximately 500 person-hours) are dedicated to preparing the Docker environment, configuring agent systems, and writing failure-triggering tests. In the future, we plan to continuously maintain and extend our benchmark to support future research on agent system maintenance. The continuous work of our benchmark is available at our website ourbenchmark . In particular, the benchmark has been extended with 20 more reproducible issues since the paper submission time. Similar trends (i.e., poor resolution rate) can be observed in those additional issues. Detailed results are presented in Appendix H.
Discussion. Based on our findings, we further discuss implications for future research towards building more effective SE agents for resolving agent issues. (i) Adding a knowledge base on agent-needed external resources. Our findings show that existing SE agents struggle with issues related to external resources. A promising direction is to augment agents with an evolving knowledge base built from API documentation, release notes, and historical issues. Integrating this knowledge could empower SE agents to better reason about and diagnose the issues related to external resources. (ii) Training SE agents with instances and trajectories collected from AgentIssue-Bench. Our benchmark and study provide training data specifically on the emerging agent systems. As our work provides executable environments and tests of buggy/fixed agent systems, future work can collect instances and trajectories (e.g., agent-environment/tool interactive trajectories) for fine-tuning more powerful SE agents that specifically targets at agent issue resolution. (iii) Adding a dynamic analysis component in SE agents. Our results highlight the limited localization accuracy of current agents, suggesting a large gap between an issue description and its root cause. To address this, future SE agent architectures could move beyond static analysis and incorporate a dynamic analysis component. By utilizing runtime information like execution trajectories and tool outputs, the agent can gather richer signals for more accurate bug localization and patch generation.
@@ 7 Conclusion
In this work, we analyze 201 GitHub issues from 18 real-world agent systems and construct the first taxonomy of agent issues. We further build AgentIssue-Bench, the first reproducible benchmark of 50 high-quality agent issue resolution tasks. Experiments on state-of-the-art SE agents demonstrate their limited effectiveness in addressing agent issues (with resolution rates ranging from 0.67% to 4.67%), highlighting the unique challenges in maintaining agent systems and the pressing need for more advanced SE agents tailored to this emerging software paradigm.
@@ Appendix A Analyzed agent systems
Agent #stars #Loc Creation time License agent-squad [4] 5.4k 109,019 07/23/2024 Apache-2.0 License AGiXT [6] 3k 111,946 04/04/2023 MIT License AI SDK [7] 18.4k 365,300 05/23/2023 Apache-2.0 License autogen [10] 44.2k 197,969 08/18/2023 CC-BY-4.0, MIT License camel [12] 12.4k 206,152 03/17/2023 Apache-2.0 License babyagi [11] 21.4k 8,800 04/03/2023 MIT License CrewAI [15] 31.3k 171,395 10/27/2023 MIT License Haystack [22] 22.9k 180,500 11/14/2019 Apache-2.0 License Lagent [24] 2.1k 13,075 08/20/2023 Apache-2.0 License MetaGPT [25] 55.4k 90,709 06/30/2023 MIT License RagaAI-Catalyst [29] 16.2k 47,252 10/01/2024 Apache-2.0 License ChatDev [13] 26.8k 40,478 11/04/2023 Apache-2.0 License gpt-engineer [20] 54.1k 17,460 04/29/2023 MIT License Pythagora [28] 1.8k 5,859 01/21/2023 Apache-2.0 License SWE-agent [31] 15.7k 63,388 04/02/2024 MIT License evo.ninja [18] 1.1k 31,862 08/18/2023 MIT License Superagent [30] 5.8k 58,602 05/10/2023 MIT License gpt-researcher [21] 21.3k 168,849 05/12/2023 Apache-2.0 License
@@ Appendix D Experiment Statistical Significance
SE Agent LLM Plausibly resolved% Correctly resolved% Localization % File-level Function-level Agentless GPT-4o 12.00 ($\pm{}4.00$) 3.33 ($\pm{}4.62$) 27.82 ($\pm{}5.51$) 12.99 ($\pm{}2.78$) Claude-3.5-S 12.00 ($\pm{}0.00$) 4.00 ($\pm{}0.00$) 27.35 ($\pm{}0.00$) 17.50 ($\pm{}0.00$) AutoCodeRover GPT-4o 7.33 ($\pm{}2.31$) 1.33 ($\pm{}2.31$) 22.07 ($\pm{}7.21$) 14.77 ($\pm{}2.62$) Claude-3.5-S 12.67 ($\pm{}6.11$) 4.67 ($\pm{}2.31$) 25.81 ($\pm{}11.43$) 19.18 ($\pm{}5.54$) SWE-agent GPT-4o 0.67 ($\pm{}2.31$) 0.67 ($\pm{}2.31$) 11.67 ($\pm{}5.17$) 4.22 ($\pm{}4.07$) Claude-3.5-S 2.00 ($\pm{}0.00$) 2.00 ($\pm{}0.00$) 9.52 ($\pm{}5.24$) 6.78 ($\pm{}2.59$)
@@ Appendix E Examples of Unresolved Issues
For the example depicted in Figure 7, the agent fails to identify the root cause of the KeyError, i.e., a conflict arising from generating multiple diffs for a single file. This issue is specific to the agent system, as it involves the handling of model outputs. However, instead of performing a deeper analysis, the agent merely prints an error message, resulting in an unsuccessful patch.
@@ F.6 Utility issues
Appendix G Comparison among SE agents with various metrics
Table 7: Overall results of SE agents on AgentIssue-Bench with task difficulty
SE Agent LLM Plausibly resolved% Correctly resolved% Localization % File-level Function-level Agentless GPT-4o 11.46 3.51 27.25 12.08 Claude-3.5-S 11.32 3.77 26.82 16.88 AutoCodeRover GPT-4o 6.84 1.40 21.62 14.37 Claude-3.5-S 11.58 4.33 25.21 18.46 SWE-agent GPT-4o 0.70 0.70 11.26 3.85 Claude-3.5-S 2.11 2.11 9.15 6.40
Table 8: Pass@1 vs. Pass@3 on AgentIssue-Bench
Pass@k GPT-4o (%) Claude-3.5-S (%) Agentless AutoCodeRover SWE-agent Agentless AutoCodeRover SWE-agent Pass@1 6.00 2.00 2.00 4.00 6.00 2.00 Pass@3 6.00 2.00 2.00 4.00 6.00 2.00
We further evaluate the fixing capabilities of SE agents by considering task difficulties. In particular, we consider the utility errors as of low severity/difficulty while the other agent-specific issues as of high severity/difficulty. Table 7 presents the weighted resolution rate (0.2 for utility errors and 0.8 for other agent-specific errors) among SE agents. Overall, we could observe similar findings between weighting and non-weighting results, including (1) limited capabilities of agents in agent issue resolution, (2) the superiority of AutoCodeRover with Claude-3.5-S, and (3) outperformance of Claude-3.5-S over GPT-4o.
Table 8 presents the pass@1 and pass@3 over one run. Overall, we could observe consistent trends on these metrics as our current findings, including (1) overall limited capabilities of agents in agent issue resolution, (2) the superiority of AutoCodeRover with Claude-3.5-S, and (3) outperformance of Claude-3.5-S over GPT-4o.
Appendix H Extended AgentIssue-Bench
Table 9 presents the results of SE agents on the 20 more issues [5] that are additionally reproduced after the paper submission time. Overall, we could observe similar trends on these additional issues that SE agents exhibit limited capabilities of resolving agent issues (i.e., up to 5% correct resolution rate).
Table 9: Results of SE agents on additional issues
SE Agent LLM Plausibly resolved% Correctly resolved% Localization % File-level Function-level Agentless GPT-4o 5.00 0.00 10.00 6.67 Claude-3.5-S 5.00 0.00 12.50 6.67 AutoCodeRover GPT-4o 5.00 0.00 12.50 5.83 Claude-3.5-S 5.00 5.00 20.83 11.25 SWE-agent GPT-4o 0.00 0.00 0.00 0.00 Claude-3.5-S 0.00 0.00 0.00 0.00